# Primitive-Root Determinant Densities over Prime Fields and Implications for PRIM-LWE

### 21 Apr 2026 v4

@@ 5.3 Pure singularity
Step 1: Tail estimate. Throughout the proof, the single prime $p=2$ is negligible in any asymptotic density statement. Fix a parameter $\delta\in(0,\tfrac{1}{2})$ and, for $y\geq 3$, set $R:=\lfloor y^{\delta}\rfloor$. Since $f$ is strongly additive, for each prime $p$ we decompose $f(p-1)=f(a_{y}(p-1))+T_{y}(p)$, where $a_{y}(p-1)$ and $T_{y}(p)$ are as in Lemma 8. By that lemma and Markov’s inequality,
$\limsup_{x\to\infty}\;\frac{1}{\pi(x)}\#\bigl\{p\leq x:T_{y}(p)>1/R^{2}\bigr\}\;\leq\;R^{2}\,\eta(y)\;\ll\;y^{2\delta-1},$ (3)
which tends to 0 as $y\to\infty$ (since $\delta<\tfrac{1}{2}$). Here and in Steps 2–3, the parameter $y$ is held fixed while $x\to\infty$; the conclusions are then strengthened by letting $y\to\infty$.
Step 2: A set of small Lebesgue measure carrying positive relative prime density. Let $m_{1}<m_{2}<\cdots<m_{R}$ be the first $R=\lfloor y^{\delta}\rfloor$ even squarefree positive integers. Since the counting function of even squarefree integers satisfies $A(t):=\#\{m\leq t:m\text{ even squarefree}\}\sim(2/\pi^{2})\,t$ (see, e.g., [39, Theorem I.3.10]), for any fixed $C>\pi^{2}/2$ we have $A(Cj)\geq j$ for all sufficiently large $j$. After enlarging $C$ if necessary, we may assume $m_{j}\leq Cj$ for all $j\geq 1$. In particular $m_{R}\leq CR$. Therefore $P^{+}(m_{j})\leq m_{j}\leq Cy^{\delta}<y$ for all sufficiently large $y$ (since $\delta<1$). Moreover,
$\sum_{j=1}^{R}\frac{1}{\varphi(m_{j})}\;\geq\;\sum_{j=1}^{R}\frac{1}{m_{j}}\;\geq\;\frac{1}{C}\sum_{j=1}^{R}\frac{1}{j}\;\gg\;\log R\;=\;\delta\log y+O(1).$ (4)
By Lemma 7, the relative density of $\{p:a_{y}(p-1)=m_{j}\}$ among the primes is $\delta_{y}(m_{j})\geq C_{y}/\varphi(m_{j})$, where $C_{y}\asymp 1/\!\log y$. Combining with (4),
$\sum_{j=1}^{R}\delta_{y}(m_{j})\;\geq\;C_{y}\sum_{j=1}^{R}\frac{1}{\varphi(m_{j})}\;\gg\;\frac{\delta\log y}{\log y}\;=\;\delta\;>\;0.$
Since the $m_{j}$ are distinct, the sets $\{p:a_{y}(p-1)=m_{j}\}$ are pairwise disjoint, so the relative density of their union $\mathcal{A}_{y}:=\{p:a_{y}(p-1)\in\{m_{1},\ldots,m_{R}\}\}$ is the sum $\sum_{j}\delta_{y}(m_{j})$. In particular, there exists an absolute constant $c_{0}>0$ (depending only on $\delta$) such that for all sufficiently large $y$, the relative density of $\mathcal{A}_{y}$ among the primes is at least $c_{0}$.
Define the closed set
$\bar{I}_{y}\;:=\;\bigcup_{j=1}^{R}\bigl[\,f(m_{j}),\;f(m_{j})+1/R^{2}\,\bigr].$
By subadditivity of Lebesgue measure, $\lambda(\bar{I}_{y})\leq R\cdot(1/R^{2})=1/R=1/\lfloor y^{\delta}\rfloor\to 0$ as $y\to\infty$. For any prime $p$ with $a_{y}(p-1)=m_{j}$ and $T_{y}(p)\leq 1/R^{2}$, the identity $f(p-1)=f(a_{y}(p-1))+T_{y}(p)$ and the nonnegativity of $T_{y}(p)$ give $f(p-1)\in[f(m_{j}),\,f(m_{j})+1/R^{2}]\subset\bar{I}_{y}$.
Step 3: Contradiction with absolute continuity. Recall that $y$ (and hence $R$) are still fixed. By (3),
$\limsup_{x\to\infty}\;\frac{1}{\pi(x)}\#\bigl\{p\leq x:T_{y}(p)>1/R^{2}\bigr\}\;=\;o_{y}(1).$
Hence, for all sufficiently large $y$,
$\liminf_{x\to\infty}\frac{1}{\pi(x)}\#\bigl\{p\leq x:f(p-1)\in\bar{I}_{y}\bigr\}\;\geq\;c_{0}-o_{y}(1)\;\geq\;\frac{c_{0}}{2}\;>\;0.$
[…]
$\mu_{f}(\bar{I}_{y})\;=\;\lim_{x\to\infty}\frac{1}{\pi(x)}\#\bigl\{p\leq x:f(p-1)\in\bar{I}_{y}\bigr\}\;\geq\;\frac{c_{0}}{2}\;>\;0.$
If $\mu_{f}$ were absolutely continuous with density $h\in L^{1}(\mathbb{R})$, then by the absolute continuity of the Lebesgue integral, $\mu_{f}(\bar{I}_{y})=\int_{\bar{I}_{y}}h\,\mathrm{d}t\to 0$ as $\lambda(\bar{I}_{y})\to 0$, contradicting $\mu_{f}(\bar{I}_{y})\geq c_{0}/2$ for all large $y$. Therefore $\mu_{f}$ is not absolutely continuous, and being of pure type and continuous, it is purely singular. ∎
Proposition 3 (Shifted-prime limiting law for $f$).
Let $f$ be as in (2). Then the empirical measures
@@
By Lemma 3, the three convergence conditions of the shifted-prime Erdős–Wintner theorem (Lemma 6) are satisfied, and the continuity criterion holds. Therefore f(p1)f(p-1) possesses a continuous limiting distribution $\mu_{f}$, which is of pure type by Lemma 6.
It remains to rule out absolute continuity. By Lemma 9, $\mu_{f}$ is purely singular. ∎
Corollary 2 (Structure of the limiting PRIM-LWE law).
The density $c(p)$ has a continuous, purely singular limiting distribution with support exactly $0,\tfrac{1}{2}]$. Its distribution function $G$ is strictly increasing on $0,\tfrac{1}{2}]$.
@@
Combine Theorems 4 and 5: $c(p)$ has a continuous limiting law with support $0,\tfrac{1}{2}]$ whose distribution function is strictly increasing. It remains to justify pure singularity.
@@
By Proposition 3, the shifted-prime law $\mu_{f}$ of f(p1)f(p-1) (with $f$ as in (2)) is continuous and purely singular. Since $\varphi(p-1)/(p-1)=e^{-f(p-1)}$ and $T(u):=e^{-u}$ is locally Lipschitz on $0,\infty)$, the push-forward $T_{\#}\mu_{f}$ preserves Lebesgue-null sets and is therefore purely singular. By Theorem 5, the same holds for $c(p)$. ∎
@@
Figure 1 compares the empirical distribution of $c(p)$ for primes $p\leq 10^{6}$ with the limiting law $G$ of Theorem 5; the close agreement already visible at this scale provides a useful numerical check on the distributional theory.
@@
Figure 1: Empirical CDF of $c(p)$ over the 78,49778{,}497 primes $3\leq p\leq 10^{6}$, overlaid with the theoretical limiting distribution function $G$ of Theorem 5, estimated by $10^{7}$ Monte Carlo samples of the infinite convolution of Lemma 10 truncated at the 10,00010{,}000th odd prime. The empirical values of $c(p)$ were computed via the truncated Euler product of Proposition 8 with $J=20$ terms, contributing error below $10^{-9}$ for every prime $p\geq 3$.
Remark 1.
@@
The value $1/2$ is not attained: $c(p)<1/2$ for every prime $p$. Indeed, for $p=2$ one has $\varphi(p-1)/(p-1)=1$, so $c(2)=\prod_{j\geq 1}(1-2^{-j})\approx 0.288788<1/2$. For odd $p$, one has $\varphi(p-1)/(p-1)\leq 1/2$ together with $\prod_{j\geq 1}(1-p^{-j})<1$. Thus $1/2$ is only a boundary point of the support of the limiting law for $c(p)$, not an attained value.
Corollary 3 (Unconditional limsup).
We have
$\limsup_{\begin{subarray}{c}p\to\infty\\ p\textup{ prime}\end{subarray}}c(p)=\frac{1}{2}.$
Moreover, for every $\varepsilon>0$, the set of primes with $c(p)>1/2-\varepsilon$ has relative density $1-G(1/2-\varepsilon)$ among the primes.
Proof.
Since $c(p)<1/2$ for every prime (Remark 1), $\limsup c(p)\leq 1/2$. For the reverse inequality, fix $\varepsilon>0$. By Corollary 2, the limiting distribution of $c(p)$ has support exactly $0,\tfrac{1}{2}]$, so the limiting measure assigns positive mass to the interval $(1/2-\varepsilon,\,1/2]$. Hence
$\lim_{x\to\infty}\frac{1}{\pi(x)}\#\bigl\{p\leq x:c(p)>1/2-\varepsilon\bigr\}=1-G(1/2-\varepsilon)>0.$
In particular, infinitely many primes satisfy $c(p)>1/2-\varepsilon$, so $\limsup c(p)\geq 1/2-\varepsilon$. Letting $\varepsilon\downarrow 0$ yields the claim. ∎
Corollary 4 (Fixed-dimension analogue).
For each fixed $n\geq 1$, the exact density $c_{n}(p)$ has the same continuous limiting distribution as $\varphi(p-1)/(p-1)$; in particular, its limiting law has support exactly $0,\tfrac{1}{2}]$. Moreover,
$\min_{\begin{subarray}{c}p\leq x\\ p\textup{ prime}\end{subarray}}c_{n}(p)\;\asymp\;\frac{1}{\log\log x}.$
Proof.
Write $c_{n}(p)=\frac{\varphi(p-1)}{p-1}P_{n}(p)$ with $P_{n}(p):=\prod_{j=1}^{n}(1-p^{-j})$. Since $P_{n}(p)\to 1$, Lemma 5 gives the distributional claim. For the lower bound on $\min_{p\leq x}c_{n}(p)$: every odd prime $p\leq x$ satisfies $\omega(p-1)\leq m(x)$, so
$\frac{\varphi(p-1)}{p-1}\;\geq\;\prod_{i=1}^{m(x)}\Bigl(1-\frac{1}{p_{i}}\Bigr),$
and $P_{n}(p)\geq 1/2$ by Lemma 1. By Lemma 2 and Mertens’ third theorem,
$\prod_{i=1}^{m(x)}\Bigl(1-\frac{1}{p_{i}}\Bigr)\;\asymp\;\frac{1}{\log\log x}.$
The upper bound is identical to Proposition 1, since $P_{n}(p)\leq 1$. ∎
6 Hausdorff Dimension, Moments, and Fourier Structure of $G$
@@
The purely singular limiting law established in Section 5 raises a natural question: how thin is the set carrying full mass of the limiting measure? Because $\mu_{f}$ decomposes as an infinite convolution of discrete measures (one for each odd prime), one can exploit the resulting independence structure to build efficient coverings and bound the Hausdorff dimension. We show that $\dim_{H}(\mu_{f})=0$, meaning that the mass concentrates on a Borel set of Hausdorff dimension zero even though the topological support is the full interval $0,\tfrac{1}{2}]$; in particular, the fractal-like clustering visible in Figure 2 is not an artifact of finite sampling. We then derive an explicit Euler product for every moment of $G$, prove that these moments determine the distribution uniquely, and establish the analytic continuation of the associated Mellin transform. The section concludes with a precise asymptotic for the behavior of $G$ near the endpoint $1/2$.
@@
Throughout, $f$ is the strongly additive function defined in (2), $\mu_{f}$ is the shifted-prime limiting measure of f(p1)f(p-1), and $X$ is a random variable with distribution $G$. By Corollary 2, $X\overset{d}{=}e^{-Y}$ where $Y\sim\mu_{f}$.
Lemma 10 (Factor structure of $\mu_{f}$).
@@
By Tenenbaum [38, formula (3), p. 262] for the shift a=1a=-1, the characteristic function of $\mu_{f}$ is
@@
For a Borel probability measure $\mu$ on $\mathbb{R}$, recall the Hausdorff dimension of $\mu$ (see, e.g., Falconer [17, Chapter 2]):
@@
Proposition 4 (Hausdorff dimension zero of the limiting law).
@@
Fix $s>0$. We shall show that $\mu_{f}$ gives full mass to a Borel set of vanishing $s$-dimensional Hausdorff measure.
For $y\geq 3$, let
@@
$\operatorname{Var}(I(y))\leq\sum_{\begin{subarray}{c}3\leq\ell\leq y\\ \ell\text{ prime}\end{subarray}}\mathbb{E}\!\left[\left(B_{\ell}\log(\ell-1)+(1-B_{\ell})\log\frac{\ell-1}{\ell-2}\right)^{2}\right]\ll\sum_{\begin{subarray}{c}3\leq\ell\leq y\\ \ell\text{ prime}\end{subarray}}\frac{(\log\ell)^{2}}{\ell}\ll(\log y)^{2}.$
Hence there exists an absolute constant $C_{0}>0$ such that
@@
Choose $C:=4C_{0}$ (so that $Cn-a_{n}s/2\leq-Cn$ in the covering estimate below).
@@
For each $n$, set $I_{n}:=I(y_{n})$ and
@@
Now define the empty-block event
@@
Mertens’ third theorem and the convergence of $\prod_{\ell\geq 3}(1-(\ell-1)^{-2})$ imply
@@
Therefore $\sum_{n}\mathbb{P}(E_{n})=\infty$. Since the random variables $B_{\ell}$ are mutually independent (Lemma 10) and the blocks $(y_{n},z_{n}]$ are pairwise disjoint, the events $E_{n}$ are mutually independent. The second Borel–Cantelli lemma therefore gives
@@
For each $\boldsymbol{b}=(b_{\ell})_{3\leq\ell\leq y_{n}}\in T_{n}$, define
@@
it follows that
@@
Because $r_{n}\downarrow 0$, for every $N\geq 1$ the set $\limsup_{n\to\infty}U_{n}$ is covered by the intervals comprising $U_{n}$ for $n\geq N$, each of diameter at most $r_{N}$. Hence
@@
Finally, $\mu_{G}$ is the push-forward of $\mu_{f}$ under the map $T(u):=e^{-u}$. Since $T$ is Lipschitz on $\log 2,\infty)$, the image of any set of Hausdorff dimension 0 again has Hausdorff dimension 0 (see, e.g., [17, Corollary 2.4]). Therefore $\dim_{H}(\mu_{G})=0$. ∎
The vanishing Hausdorff dimension of the limiting measure explains the clustering visible in Figure 2: the purely singular law forces the values of $c(p)$ to concentrate near arithmetic anchor points of the form $\frac{1}{2}\prod_{r\in S}(1-1/r)$, determined by the small prime divisors of p1p-1, rather than spreading uniformly over $0,\frac{1}{2}]$.
Figure 2: Empirical bin masses of $c(p)$ over the 78,49778{,}497 primes $3\leq p\leq 10^{6}$, computed via Proposition 8 with truncation depth $J=20$, using $100$ equal-width bins covering $[0,0.5]$. The dashed bars mark selected anchor values $\varphi(m)/m$ for small even squarefree $m$. The concentration of mass near these arithmetic anchors reflects the dependence of $c(p)$ on the small prime divisors of p1p-1 and illustrates the clustering predicted by the purely singular limiting law of Corollary 2.
Theorem 6 (Moment Euler product).

### 06 May 2026 v5

@@ 5.3. Pure singularity
Step 1: Tail estimate. Fix $\delta\in(0,\tfrac{1}{2})$ and set $R:=\lfloor y^{\delta}\rfloor$. Decompose $f(p-1)=f(a_{y}(p-1))+T_{y}(p)$. By Lemma 5.11 and Markov’s inequality,
(5.4) $\limsup_{x\to\infty}\;\frac{1}{\pi(x)}\#\bigl\{p\leq x:T_{y}(p)>1/R^{2}\bigr\}\;\leq\;R^{2}\,\eta(y)\;\ll\;y^{2\delta-1}\;\xrightarrow{y\to\infty}\;0.$
Step 2: Small-measure set with positive prime density. Let $m_{1}<\cdots<m_{R}$ be the smallest $R$ even squarefree positive integers. Since even squarefree integers have positive asymptotic density (equal to $2/\pi^{2}$), there is an absolute constant $C>0$ with $m_{j}\leq Cj$; in particular $m_{j}\leq CR=C\lfloor y^{\delta}\rfloor\ll y^{1/2}$ since $\delta<\tfrac{1}{2}$, so $P^{+}(m_{j})\leq m_{j}<y$ for all sufficiently large $y$. Because $\varphi(m_{j})\leq m_{j}\leq Cj$, we have $\sum_{j\leq R}1/\varphi(m_{j})\geq C^{-1}\sum_{j\leq R}1/j\gg\log R$. By Lemma 5.10, $\delta_{y}(m_{j})\geq Q(y)/\varphi(m_{j})$ with $Q(y)\asymp 1/\!\log y$, so
(5.5) $\sum_{j=1}^{R}\delta_{y}(m_{j})\;\geq\;Q(y)\sum_{j=1}^{R}\frac{1}{\varphi(m_{j})}\;\gg\;\frac{\log R}{\log y}\;=\;\delta+o(1)\;\gg_{\delta}\;1$
for all sufficiently large $y$, where the implied constants are absolute (from Mertens’ theorem and the density of even squarefree integers) and $\log R/\!\log y=\delta+o(1)$ because $R=\lfloor y^{\delta}\rfloor$. Define the closed set $\bar{I}_{y}:=\bigcup_{j=1}^{R}[f(m_{j}),\,f(m_{j})+1/R^{2}]$. Then $\lambda(\bar{I}_{y})\leq 1/R\to 0$ as $y\to\infty$, while for any $p$ with $a_{y}(p-1)=m_{j}$ and $T_{y}(p)\leq 1/R^{2}$ one has $f(p-1)\in\bar{I}_{y}$.
Step 3: Contradiction. Since $a_{y}(p-1)$ takes a unique value for each prime $p$, the sets $\{p:a_{y}(p-1)=m_{j}\}$ are pairwise disjoint. Finite additivity of the natural density therefore gives $\lim_{x\to\infty}\frac{1}{\pi(x)}\#\{p\leq x:a_{y}(p-1)\in\{m_{1},\dots,m_{R}\}\}=\sum_{j\leq R}\delta_{y}(m_{j})$. Removing the primes for which $T_{y}(p)>1/R^{2}$ loses at most the quantity in (5.4), so combining with (5.5): there exists $c_{0}>0$ independent of $y$ such that for all large $y$,
$\mu_{f}(\bar{I}_{y})\;=\;\lim_{x\to\infty}\frac{1}{\pi(x)}\#\bigl\{p\leq x:f(p-1)\in\bar{I}_{y}\bigr\}\;\geq\;c_{0}\;>\;0,$
the limit being justified by the Portmanteau theorem (see, e.g., [2, Theorem 2.1]): $\bar{I}_{y}$ is a $\mu_{f}$-continuity set because its boundary consists of at most $2R$ points and $\mu_{f}$ is atom-free (Theorem 5.9(b)), so weak convergence of the empirical measures to $\mu_{f}$ gives $\mu_{x}(\bar{I}_{y})\to\mu_{f}(\bar{I}_{y})$. Since $\lambda(\bar{I}_{y})\to 0$ while $\mu_{f}(\bar{I}_{y})\geq c_{0}$, the measure $\mu_{f}$ cannot be absolutely continuous. Being continuous and of pure type, it is purely singular. ∎
Proposition 5.13 (Shifted-prime limiting law for $f$).
Let $f$ be as in (5.1). Then the empirical measures
@@
By Lemma 5.1, the three convergence conditions of the shifted-prime Erdős–Wintner theorem (Theorem 5.9) are satisfied, and the continuity criterion holds. Therefore f(p1)f(p-1) possesses a continuous limiting distribution $\mu_{f}$, which is of pure type by Theorem 5.9.
It remains to rule out absolute continuity. By Lemma 5.12, $\mu_{f}$ is purely singular. ∎
Corollary 5.14 (Structure of the limiting law for $c(p)$).
The ratio $c(p)$ has a continuous, purely singular limiting distribution with support exactly $0,\tfrac{1}{2}]$. Its distribution function $G$ is strictly increasing on $0,\tfrac{1}{2}]$.
@@
Combine Theorems 5.5 and 5.8: $c(p)$ has a continuous limiting law with support $0,\tfrac{1}{2}]$ whose distribution function is strictly increasing. It remains to justify pure singularity.
@@
By Proposition 5.13, the shifted-prime law $\mu_{f}$ of f(p1)f(p-1) (with $f$ as in (5.1)) is continuous and purely singular. Since $\varphi(p-1)/(p-1)=e^{-f(p-1)}$ and $T(u):=e^{-u}$ is Lipschitz on $0,\infty)$ (since $|T^{\prime}(u)|=e^{-u}\leq 1$), the push-forward $T_{\#}\mu_{f}$ preserves Lebesgue-null sets. Moreover, $T$ is injective on $\log 2,\infty)$ and $\mu_{f}$ concentrates on this half-line (since $Y\geq\log 2$ a.s.), so the carrying null set of $\mu_{f}$ maps bijectively to a null set carrying $T_{\#}\mu_{f}$. Hence $T_{\#}\mu_{f}$ is purely singular. Since Theorem 5.8 shows that $c(p)$ has the same limiting distribution $G$ as $\varphi(p-1)/(p-1)$, the limiting measure for $c(p)$ is exactly the same measure $\mu_{G}$, and pure singularity carries over. ∎
The strictly stronger metric result $\dim_{H}(\mu_{G})=0$ is established in Theorem 6.2 by an independent method (a Borel–Cantelli covering argument exploiting the Bernoulli-convolution independence structure).
@@
Figure 1 compares the empirical distribution of $c(p)$ for primes $p\leq 10^{6}$ with the limiting law $G$ of Theorem 5.8; the close agreement already visible at this scale provides a useful numerical check on the distributional theory.
Remark 5.15 (Truncation error for numerical approximation).
Truncate the Bernoulli series of Lemma 6.1 at the first $N$ odd primes ($\ell\leq\ell_{N}$) and let $G_{N}$ denote the distribution function of the truncated law $X_{N}:=\tfrac{1}{2}\prod_{3\leq\ell\leq\ell_{N}}(1-1/\ell)^{B_{\ell}}$. Since $|X-X_{N}|\leq\tfrac{1}{2}T_{N}$ almost surely, where $T_{N}:=\sum_{\ell>\ell_{N},\,\ell\textup{ prime}}B_{\ell}\log(\ell/(\ell-1))$, the coupling characterisation of the Wasserstein-1 distance gives
$W_{1}(G_{N},G)\;\leq\;\tfrac{1}{2}\,\mathbb{E}[T_{N}]\;=\;\tfrac{1}{2}\sum_{\begin{subarray}{c}\ell>\ell_{N}\\ \ell\textup{ prime}\end{subarray}}\frac{1}{\ell-1}\log\frac{\ell}{\ell-1}\,.$
For N=10, [^0] N=10{,}000 (so $\ell_{N}=104{,}743$), $\mathbb{E}[T_{N}]<10^{-5}$ and therefore $W_{1}(G_{N},G)<5\times 10^{-6}$. Since $G_{N}$ has an atom of mass $Q(\ell_{N})\approx 0.064$ at $\tfrac{1}{2}$—where $Q(y)$ is the product defined in Lemma 5.10, equal to the probability that every Bernoulli indicator up to $\ell_{N}$ equals zero—while $G$ is continuous, the truncation error is controlled in $W_{1}$, not in the Kolmogorov metric. The plotted theoretical curve is a Monte Carlo approximation to $G_{N}$; the DKW inequality controls the sampling error (see Figure 1).
@@
Figure 1. Empirical CDF of $c(p)$ over the 78,49778{,}497 primes $3\leq p\leq 10^{6}$, overlaid with the theoretical CDF $G$ of Theorem 5.8, approximated by $10^{7}$ Monte Carlo samples of the truncated convolution $G_{N}$ with N=10, [^0] N=10{,}000 odd primes. See Remark 5.15 for truncation and sampling error bounds. Empirical $c(p)$ values use Remark 7.9 with $J=20$.
Remark 5.16 (The value $1/2$: boundary but not attained).
@@
The value $1/2$ is not attained: $c(p)<1/2$ for every prime $p$. Indeed, for $p=2$ one has $\varphi(p-1)/(p-1)=1$, so $c(2)=\prod_{j\geq 1}(1-2^{-j})\approx 0.2888<1/2$. For odd $p$, one has $\varphi(p-1)/(p-1)\leq 1/2$ together with $\prod_{j\geq 1}(1-p^{-j})<1$. Thus $1/2$ is only a boundary point of the support, not an attained value. Since the support of $G$ is exactly $0,\tfrac{1}{2}]$ (Corollary 5.14), it follows that $\limsup_{p\to\infty,\,p\textup{ prime}}c(p)=1/2$; for every $\varepsilon>0$, the set of primes with $c(p)>1/2-\varepsilon$ has relative density $1-G(1/2-\varepsilon)>0$ among the primes.
Remark 5.17 (Fixed-dimension analogue).
For each fixed $n\geq 1$, $c_{n}(p)=\frac{\varphi(p-1)}{p-1}P_{n}(p)$ with $P_{n}(p):=\prod_{j=1}^{n}(1-p^{-j})\to 1$, so Lemma 5.7 gives $c_{n}(p)$ the same limiting distribution as $\varphi(p-1)/(p-1)$, and the primorial/Mertens argument yields $\min_{p\leq x,\,p\textup{ prime}}c_{n}(p)\asymp 1/\!\log\log x$.
6. Hausdorff Dimension, Moments, and Fourier Structure
@@
The purely singular limiting law established in Section 5 raises a natural question: how thin is the set carrying full mass of the limiting measure? Theorem 6.2 below provides the strictly stronger result $\dim_{H}(\mu_{f})=0$, giving in particular an independent proof of pure singularity. The Erdős-style singularity argument in Section 5.3 is included for its independent methodological interest. Because $\mu_{f}$ decomposes as an infinite convolution of discrete measures (one for each odd prime), one can exploit the resulting independence structure to build efficient coverings and bound the Hausdorff dimension. We show that $\dim_{H}(\mu_{f})=0$, meaning that the mass concentrates on a Borel set of Hausdorff dimension zero even though the topological support of the push-forward $\mu_{G}$ is the full interval $0,\tfrac{1}{2}]$. We then derive an explicit Euler product for every moment of $G$, prove that these moments determine the distribution uniquely, and establish the analytic continuation of the associated Mellin transform. The section concludes with a precise asymptotic for the behaviour of $G$ near the endpoint $1/2$.
@@
Throughout, $f$ is the strongly additive function defined in (5.1), $\mu_{f}$ is the shifted-prime limiting measure of f(p1)f(p-1), and $X$ is a random variable with distribution $G$. By Corollary 5.14, $X\overset{d}{=}e^{-Y}$ where $Y\sim\mu_{f}$. The Bernoulli indicators $B_{\ell}$ and the scaled jumps $Z_{\ell}:=B_{\ell}\,w_{\ell}$ (with $w_{\ell}=\log(\ell/(\ell-1))$) are used interchangeably below; the former appears in the covering construction (Theorem 6.2), the latter in the convolution decomposition (Lemma 6.1).
6.1. Convolution structure and Hausdorff dimension
Lemma 6.1 (Factor structure of $\mu_{f}$).
@@ 6.1. Convolution structure and Hausdorff dimension
By Theorem 5.9, the characteristic function of $\mu_{f}$ is given by (5.3):
@@
For a Borel probability measure $\mu$ on $\mathbb{R}$, recall the Hausdorff dimension of $\mu$ (see, e.g., Falconer [14, Chapter 2]):
@@
Theorem 6.2 (Hausdorff dimension zero of the limiting law).
@@
Fix $s>0$. We shall show that $\mu_{f}$ gives full mass to a Borel set of vanishing $s$-dimensional Hausdorff measure. The argument proceeds in five steps: (1) construct recursively a sequence of super-exponentially growing covering scales; (2) bound the number of prefix configurations via Chebyshev’s inequality; (3) show that empty-block events occur infinitely often by the second Borel–Cantelli lemma; (4) control the tail contribution by Markov’s inequality; (5) assemble the covering and estimate its Hausdorff measure.
Step 1: Recursive covering scales. For $y\geq 3$, let
@@
$\operatorname{Var}(I(y))$ $\leq\sum_{\begin{subarray}{c}3\leq\ell\leq y\\ \ell\text{ prime}\end{subarray}}\mathbb{E}\!\left[\left(B_{\ell}\log(\ell-1)+(1-B_{\ell})\log\frac{\ell-1}{\ell-2}\right)^{\!2}\right]$
$\ll\sum_{\begin{subarray}{c}3\leq\ell\leq y\\ \ell\text{ prime}\end{subarray}}\frac{(\log\ell)^{2}}{\ell}\;\ll\;(\log y)^{2},$
where the final estimate follows by partial summation and the prime number theorem. Hence there exists an absolute constant $C_{0}>0$ such that
@@
Choose $C:=4C_{0}$ and define
@@
Since $a_{n}\geq 4Cn/s$, it follows that
[…]
this inequality drives the convergence of the covering estimate below.
Step 2: Concentration and cardinality of prefix configurations. For each $n$, set $I_{n}:=I(y_{n})$ and
@@
Step 3: Empty-block divergence (second Borel–Cantelli). Define the empty-block event
@@
Mertens’ third theorem and the convergence of $\prod_{\ell\geq 3}(1-(\ell-1)^{-2})$ imply that the product of $(1-1/\ell)$ over primes in $(y_{n},z_{n}]$ is asymptotic to $(\log y_{n})/(\log z_{n})=1/a_{n}$, since $\log z_{n}=a_{n}\log y_{n}$. Therefore
@@
Therefore $\sum_{n}\mathbb{P}(E_{n})=\infty$. Since the random variables $B_{\ell}$ are mutually independent (Lemma 6.1) and the blocks $(y_{n},z_{n}]$ are pairwise disjoint, the events $E_{n}$ are mutually independent. The second Borel–Cantelli lemma therefore gives
@@
Step 5: Covering and Hausdorff measure estimate. For each $\boldsymbol{b}=(b_{\ell})_{3\leq\ell\leq y_{n}}\in T_{n}$, define
@@
the first Borel–Cantelli lemma implies that $A_{n}$ and $F_{n}$ each hold for all sufficiently large $n$ almost surely, while the second Borel–Cantelli lemma (applicable because the $E_{n}$ are independent) gives $E_{n}$ infinitely often almost surely. Hence $A_{n}\cap E_{n}\cap F_{n}$ occurs infinitely often almost surely, and on each such occurrence $Y\in U_{n}$. Therefore
@@
Because $r_{n}\downarrow 0$, for every $N\geq 1$ the set $\limsup_{n\to\infty}U_{n}$ is covered by the intervals comprising $U_{n}$ for $n\geq N$, each of diameter at most $r_{N}$. Denoting by $\mathcal{H}^{s}_{\rho}(E):=\inf\bigl\{\sum_{i}(\operatorname{diam}E_{i})^{s}:E\subset\bigcup_{i}E_{i},\;\operatorname{diam}E_{i}\leq\rho\bigr\}$ the $s$-dimensional Hausdorff pre-measure at scale $\rho$ (see [14, Section 2.1]),
@@
Finally, $\mu_{G}$ is the push-forward of $\mu_{f}$ under the map $T(u):=e^{-u}$. Since every $Z_{\ell}\geq 0$, we have $Y\geq\log 2$ almost surely, so the carrying Borel set lies in $\log 2,\infty)$. Since $T$ is Lipschitz on $\log 2,\infty)$ (with $|T^{\prime}(u)|=e^{-u}\leq\tfrac{1}{2}$), the image of any set of Hausdorff dimension 0 again has Hausdorff dimension 0 (see, e.g., [14, Corollary 2.4]). Therefore $\dim_{H}(\mu_{G})=0$. ∎
The vanishing Hausdorff dimension of the limiting measure explains why the values of $c(p)$ concentrate near arithmetic anchor points of the form $\frac{1}{2}\prod_{r\in S}(1-1/r)$, determined by the small prime divisors of p1p-1, rather than spreading uniformly over $0,\frac{1}{2}]$.
Figure 2. Empirical bin masses of $c(p)$ over the 78,49778{,}497 primes $3\leq p\leq 10^{6}$, computed via Remark 7.9 with truncation depth $J=20$, using $100$ equal-width bins covering $[0,0.5]$. The dashed bars mark selected anchor values $\varphi(m)/m$ for small even squarefree $m$. The concentration of mass near these arithmetic anchors reflects the dependence of $c(p)$ on the small prime divisors of p1p-1 and illustrates the clustering predicted by the purely singular limiting law of Corollary 5.14.
6.2. Moments and Mellin transform
Theorem 6.3 (Moment Euler product).